OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region underneath the graph of f(x)=(x)/(sqrt(x^3+5)), about the y-axis over the interval [1, 10].

OpenStudy (anonymous):

which method do you want to use to do this?

OpenStudy (anonymous):

using the shell method you would integrate from a to be over 2 pi (shell radius) (shell height) dx.

OpenStudy (anonymous):

\[\int\limits_{1}^{10}2\pi*(x)(\frac{ x }{ x^3 }+5)\]

OpenStudy (anonymous):

pull the 2 pi out front and distribute the x. then simplify

OpenStudy (anonymous):

i can do it from here

OpenStudy (anonymous):

woo okay. good luck!

OpenStudy (anonymous):

i integrated it and got 1570

OpenStudy (anonymous):

\[\int\limits\limits_{1}^{10}2\pi*(x)\frac{ x }{ \sqrt{x^3+5} }dx\]

OpenStudy (anonymous):

yeah?

OpenStudy (anonymous):

\[2\pi \int\limits_{1}^{10}\frac{ x^2 }{(x^3+5)^{1/2} }dx\]

OpenStudy (anonymous):

this looks nasty so lets look for a substitution. let u = x^3+5... du = 3x^2 dx...1/3 du = x^2 dx

OpenStudy (anonymous):

your answer was not correct.. why i am still explaining.

OpenStudy (anonymous):

your answer is only of by 1400+

OpenStudy (anonymous):

answer is 122, i see what i did wrong.